# Blade Antenna-SDR System Prototype for the CANTAR Global 21-cm Experiment: Simulations, Measurements, and In-Situ Results

### v1

@@
Abstract
We present the design and initial testing of a low-frequency radio telescope prototype developed for the CANTAR (Colombian Antarctic Telescope for 21-cm Absorption during Reionization) experiment. Operating from 100 to 200 MHz, the system integrates a blade dipole antenna inspired by the EDGES high-band design with a software-defined radio (SDR) receiver. We report simulations of antenna impedance and beam chromaticity, along with SDR performance tests (Limenet Mini, Ettus E310, USRP2920). A dual-stage low-noise amplifier reduces system temperature, enabling foreground-sensitive observations. Radiometric estimates suggest sub-mK sensitivity is achievable with 1000 h of integration. This prototype forms part of Colombia’s emerging infrastructure for 21-cm cosmology, with deployments planned in low-RFI sites in the Colombian Andes and Antarctica.
@@ 0.1 Introduction
The redshifted 21-cm line of neutral hydrogen provides a unique observational window into the thermal and ionization history of the early universe. In particular, measurements of the global (sky-averaged) 21-cm signal can probe the Cosmic Dawn and the Epoch of Reionization (EoR) (Barkana & Loeb, 2005; Furlanetto et al., 2006; Mozdzen et al., 2016). During this period, the first luminous sources coupled the hydrogen spin temperature to the kinetic temperature of the gas via the Wouthuysen-Field effect, producing an absorption signature in the 21-cm line against the cosmic microwave background (CMB) (Costa et al., 2018). The frequency range corresponding to this evolution lies between approximately 40 and 200 MHz for redshifts $6\leq z\leq 30$ (Bharadwaj & Ali, 2004), with an amplitude of $\sim$100–200 mK and width of $\sim$10 MHz, depending on the astrophysical heating history.
This signal is embedded in a sky dominated by bright galactic synchrotron emission and extragalactic foregrounds that are several orders of magnitude stronger ($\sim 10^{3}-10^{4}$ K). Its detection demands tight control over instrumental systematics, including beam chromaticity, impedance mismatches, and thermal gain variations (Bowman et al., 2009; Bernardi et al., 2016; Monsalve et al., 2017; Cheng et al., 2018; Restrepo et al., 2023). The controversial EDGES detection of an absorption feature at 78 MHz (Bowman et al., 2018), with an amplitude exceeding theoretical expectations, has renewed interest in independent validation and motivated a new generation of experiments (Monsalve et al., 2024).
Most global 21-cm efforts use single-element radiometers with fixed analog backends. These include EDGES, SARAS (Singh et al., 2018), PRIZM (Philip et al., 2019), REACH (de Lera Acedo et al., 2022), BIGHORNS (Sokolowski et al., 2015). Though differing in calibration approach, antenna type, and beam chromaticity control, they share the goal of suppressing spectral structure in the instrument response. Interferometric arrays such as HERA (DeBoer et al., 2017), LEDA (Price et al., 2018), and LOFAR (Ghara et al., 2025) pursue complementary spatially resolved 21-cm measurements.
@@
Unlike purpose-built analog backends, SDRs provide flexibility and rapid prototyping capabilities in radio science applications, but raise concerns about spectral fidelity due to ADC resolution, phase noise, clock instability, and internal filtering (Feng et al., 2021; Ragoomundun & Beeharry, 2019; Keller et al., 2015). One aim of this work is to quantify these limitations and assess whether SDRs can serve as viable alternatives for radiometer-based 21-cm cosmology when combined with analog filtering and appropriate calibration strategies.
@@ 0.2.1 System Requirements
The design of the prototype system presented here is guided by constraints common to global 21-cm signal experiments like suppressing spectral structure in the system response and achieving wideband impedance matching. While the antenna does not require high spatial resolution, it must provide maximum gain at zenith and maintain a return loss better than 10 dB, gain between 10-30 dB, and a noise figure below 4 dB (Restrepo et al., 2023). Table 1 summarizes the relevant characteristics of comparable global signal experiments.
@@ 0.2.2 Role of SDR Technology
Software-defined radios (SDRs) provide a reconfigurable, low-cost backend with tunable sampling rates, bandwidths, and gain. Despite their widespread use in RFI monitoring and radiometry (Keller et al., 2015; Vacaliuc et al., 2011), SDRs remain underused in global 21-cm experiments due to their limited dynamic range, internal filtering stages, and susceptibility to quantization and clock drift artifacts.
@@ 0.2.3 Radiotelescope System Description
A Roberts (or Collins) balun (Rogers, 2012) connects the balanced dipole to the unbalanced coaxial transmission line. This balun consists of two copper tubes and a sliding tuning plate, which allows for mechanical adjustment of the electrical length and impedance match. By varying the tuner height, the resonant frequency and bandwidth of the antenna can be shifted. A metal enclosure was added to suppress common-mode currents along the feed structure. The output port is a standard SMA connector located at the balun base.
@@ 0.3.2 Beam Chromaticity
In both planes, the antenna exhibits reduced chromatic variation at zenith and increased variation near the horizon, as expected. For $\phi=0^{\circ}$, the gain variation below 20 elevation exceeds 3 dB over the band, while at higher elevations it remains relatively flat. For $\phi=90^{\circ}$, the behavior is broadly similar, with the smoothest region occurring between 60 and zenith. These results are consistent with expectations for a blade dipole and suggest that the antenna is sufficiently stable for beam modeling and calibration, especially if zenith-pointing observations are prioritized (Restrepo et al., 2023).
@@ 0.3.4 Antenna performance metrics
$A_{\mathrm{e}}=\frac{\lambda^{2}G}{4\pi}=\frac{(2\,\mathrm{m})^{2}\cdot 5.01}{4\pi}\approx 1.59\leavevmode\nobreak\ \mathrm{m}^{2},$ (1)
@@
$T_{\mathrm{src}}=\frac{A_{\mathrm{e}}}{2k}\cdot S=\frac{1.59}{2k}\cdot 1.24\times 10^{-20}\leavevmode\nobreak\ \mathrm{W\,m^{-2}\,Hz^{-1}}\approx 716\leavevmode\nobreak\ \mathrm{K}.$ (2)
@@
$T_{\mathrm{sys}}=T_{\mathrm{ant}}+T_{\mathrm{rx}}\approx 300\leavevmode\nobreak\ \mathrm{K}+1370\leavevmode\nobreak\ \mathrm{K}=1670\leavevmode\nobreak\ \mathrm{K},$ (3)
@@
$T_{\mathrm{sys}}\approx 300\leavevmode\nobreak\ \mathrm{K}+153\leavevmode\nobreak\ \mathrm{K}=453\leavevmode\nobreak\ \mathrm{K},$ (4)
@@
The proposed configuration adds high- and low-pass filters to suppress out-of-band power and improve gain flatness. Two ZX60-33LNR-S+ amplifiers, each with 24.7 dB gain and 1.1 dB NF, are combined with Mini-Circuits SHP-20+ and SLP-250+ filters. These filters restrict the frequencies adjacent to those of interest (between 100 and 200 MHz) (see Figure 13). This architecture is based on the pre-amplification stage used in the MIST and MINI-MIST projects (Restrepo et al., 2020). The resulting system temperature is further reduced:
$T_{\mathrm{sys}}\approx 300\leavevmode\nobreak\ \mathrm{K}+143\leavevmode\nobreak\ \mathrm{K}=443\leavevmode\nobreak\ \mathrm{K},$ (5)
@@ 0.4.2 Radiometric Sensitivity and Detection Thresholds
$\Delta T=\frac{443}{\sqrt{10^{6}\times 3.6\times 10^{6}}}\approx 0.23\leavevmode\nobreak\ \mathrm{mK}.$ (7)
This result indicates that, under ideal observing conditions and assuming perfect calibration and system stability, the instrument is thermally capable of detecting signals at the sub-millikelvin level. Since the anticipated global 21-cm signal during the Cosmic Dawn has an amplitude on the order of 100 mK, the current system—at least from a thermal noise perspective—is sufficiently sensitive. For comparison, the HERA experiment used a similar integration time of 94 nights to place stringent upper limits on the 21 cm power spectrum during reionization, reaching sensitivities sufficient to rule out a broad class of cold IGM models (Facchinetti et al., 2024).
@@ 0.5 Future Work
We are developing novel, software optimized antenna designs (Restrepo et al., 2023) along with radio sky model validation of antenna chromaticity and bayesian foreground models (Mora et al, in prep).
@@ 0.6 Conclusions
Compared to established radiometers such as EDGES, SARAS, PRIZM, and BIGHORNS, this CANTAR prototype demonstrates broadly consistent performance in key areas such as impedance match and system temperature, while introducing a novel digital backend architecture based on commercial software-defined radios (SDRs). For instance, our best-performing antenna configurations achieve simulated and measured $S_{11}$ values below –25 dB over bandwidths of 50–70 MHz, which is comparable to the –15 to –20 dB targets reported for EDGES high-band (Bowman et al., 2009) and SARAS 3 (Singh et al., 2018). Similarly, the prototype system temperature of 443–453 K using a dual-stage low-noise amplifier is within the 300–500 K range reported for BIGHORNS and PRIZM (Sokolowski et al., 2015; Philip et al., 2019).

### v2

@@
Blade Antenna-SDR System Prototype for the CANTAR Global 21-cm Experiment: Simulations, Measurements, and In-Situ Results
Felipe P. Mosquera11affiliation: Escuela de Ingenierías Eléctrica, Electrónica y Telecomunicaciones. Universidad Industrial de Santander,Universidad Industrial de Santander, Bucaramanga, Colombia. 44affiliation: Fundación para el Desarrollo de la Radioastronomía y Tecnologías Aplicadas - FUDARTA, Bogotá, Colombia.    Julian Rodriguez-Ferreira11affiliationmark:    Efrén Acevedo11affiliationmark:
Oscar Restrepo22affiliation: Universidad ECCI, Bogotá, Colombia. 44affiliationmark:
   David González11affiliationmark:    & Germán Chaparro33affiliation: Grupo de Física y Astrofísica Computacional, FACOM, Universidad de Antioquia, Medellín, Colombia. 44affiliationmark: Efren Acevedo, David Gonzalez, Felipe P. Mosquera, Julian Rodriguez-Ferreira: Universidad Industrial de Santander, 680002, Carrera 27 Calle 9, Bucaramanga, Santander, Colombia (efrenacevedoc@gmail.com, david2248098@correo.uis.edu.co, felipe.mosquera@fudarta.org, jgrodrif@uis.edu.co). German Chaparro: Universidad de Antioquia, Calle 67 N. 53-108, 050010, Medellin, Antioquia, Colombia (german.chaparro@udea.edu.co). Oscar Restrepo: Universidad ECCI, Carrera 19 N. 49 - 20, 110221, Bogotá, Cundinamarca, Colombia (orestrepog@ecci.edu.co).
@@ 0.1 Introduction
The redshifted 21-cm line of neutral hydrogen provides a unique observational window into the thermal and ionization history of the early universe. In particular, measurements of the global (sky-averaged) 21-cm signal can probe the Cosmic Dawn and the Epoch of Reionization (EoR) (barkana; Furlanetto_2006; 2016MNRAS.455.3890M). During this period, the first luminous sources coupled the hydrogen spin temperature to the kinetic temperature of the gas via the Wouthuysen-Field effect, producing an absorption signature in the 21-cm line against the cosmic microwave background (CMB) (costa). The frequency range corresponding to this evolution lies between approximately 40 and 200 MHz for redshifts $6\leq z\leq 30$ (article6), with an amplitude of $\sim$100–200 mK and width of $\sim$10 MHz, depending on the astrophysical heating history.
This signal is embedded in a sky dominated by bright galactic synchrotron emission and extragalactic foregrounds that are several orders of magnitude stronger ($\sim 10^{3}-10^{4}$ K). Its detection demands tight control over instrumental systematics, including beam chromaticity, impedance mismatches, and thermal gain variations (bowman2009; bernardi; Monsalve_2017; Cheng; restrepo). The controversial EDGES detection of an absorption feature at 78 MHz (2bc8363417fb44159194d3775d6e51bd), with an amplitude exceeding theoretical expectations, has renewed interest in independent validation and motivated a new generation of experiments (mist2023).
Most global 21-cm efforts use single-element radiometers with fixed analog backends. These include EDGES, SARAS (Singh_2018), PRIZM (prizm), REACH (reach), BIGHORNS (Sokolowski_2015). Though differing in calibration approach, antenna type, and beam chromaticity control, they share the goal of suppressing spectral structure in the instrument response. Interferometric arrays such as HERA (DeBoer_2017), LEDA (10.1093), and LOFAR (6051249) pursue complementary spatially resolved 21-cm measurements.
@@
Unlike purpose-built analog backends, SDRs provide flexibility and rapid prototyping capabilities in radio science applications, but raise concerns about spectral fidelity due to ADC resolution, phase noise, clock instability, and internal filtering (article1; sdrragoo; 7303211). One aim of this work is to quantify these limitations and assess whether SDRs can serve as viable alternatives for radiometer-based 21-cm cosmology when combined with analog filtering and appropriate calibration strategies.
@@ 0.2.1 System Requirements
The design of the prototype system presented here is guided by constraints common to global 21-cm signal experiments like suppressing spectral structure in the system response and achieving wideband impedance matching. While the antenna does not require high spatial resolution, it must provide maximum gain at zenith and maintain a return loss better than 10 dB, gain between 10-30 dB, and a noise figure below 4 dB (restrepo). Table 1 summarizes the relevant characteristics of comparable global signal experiments.
@@ 0.2.2 Role of SDR Technology
Software-defined radios (SDRs) provide a reconfigurable, low-cost backend with tunable sampling rates, bandwidths, and gain. Despite their widespread use in RFI monitoring and radiometry (7303211; osti_1074425), SDRs remain underused in global 21-cm experiments due to their limited dynamic range, internal filtering stages, and susceptibility to quantization and clock drift artifacts.
@@ 0.2.3 Radiotelescope System Description
A Roberts (or Collins) balun (inbook1100) connects the balanced dipole to the unbalanced coaxial transmission line. This balun consists of two copper tubes and a sliding tuning plate, which allows for mechanical adjustment of the electrical length and impedance match. By varying the tuner height, the resonant frequency and bandwidth of the antenna can be shifted. A metal enclosure was added to suppress common-mode currents along the feed structure. The output port is a standard SMA connector located at the balun base.
@@ 0.3.2 Beam Chromaticity
In both planes, the antenna exhibits reduced chromatic variation at zenith and increased variation near the horizon, as expected. For $\phi=0^{\circ}$, the gain variation below 20 elevation exceeds 3 dB over the band, while at higher elevations it remains relatively flat. For $\phi=90^{\circ}$, the behavior is broadly similar, with the smoothest region occurring between 60 and zenith. These results are consistent with expectations for a blade dipole and suggest that the antenna is sufficiently stable for beam modeling and calibration, especially if zenith-pointing observations are prioritized (restrepo).
@@ 0.3.4 Antenna performance metrics
$A_{\mathrm{e}}=\frac{\lambda^{2}G}{4\pi}=\frac{(2\,\mathrm{m})^{2}\cdot 5.01}{4\pi}\approx 1.59\penalty 10000\ \mathrm{m}^{2},$ (1)
@@
$T_{\mathrm{src}}=\frac{A_{\mathrm{e}}}{2k}\cdot S=\frac{1.59}{2k}\cdot 1.24\times 10^{-20}\penalty 10000\ \mathrm{W\,m^{-2}\,Hz^{-1}}\approx 716\penalty 10000\ \mathrm{K}.$ (2)
@@
$T_{\mathrm{sys}}=T_{\mathrm{ant}}+T_{\mathrm{rx}}\approx 300\penalty 10000\ \mathrm{K}+1370\penalty 10000\ \mathrm{K}=1670\penalty 10000\ \mathrm{K},$ (3)
@@
$T_{\mathrm{sys}}\approx 300\penalty 10000\ \mathrm{K}+153\penalty 10000\ \mathrm{K}=453\penalty 10000\ \mathrm{K},$ (4)
@@
The proposed configuration adds high- and low-pass filters to suppress out-of-band power and improve gain flatness. Two ZX60-33LNR-S+ amplifiers, each with 24.7 dB gain and 1.1 dB NF, are combined with Mini-Circuits SHP-20+ and SLP-250+ filters. These filters restrict the frequencies adjacent to those of interest (between 100 and 200 MHz) (see Figure 13). This architecture is based on the pre-amplification stage used in the MIST and MINI-MIST projects (rest). The resulting system temperature is further reduced:
$T_{\mathrm{sys}}\approx 300\penalty 10000\ \mathrm{K}+143\penalty 10000\ \mathrm{K}=443\penalty 10000\ \mathrm{K},$ (5)
@@ 0.4.2 Radiometric Sensitivity and Detection Thresholds
$\Delta T=\frac{443}{\sqrt{10^{6}\times 3.6\times 10^{6}}}\approx 0.23\penalty 10000\ \mathrm{mK}.$ (7)
This result indicates that, under ideal observing conditions and assuming perfect calibration and system stability, the instrument is thermally capable of detecting signals at the sub-millikelvin level. Since the anticipated global 21-cm signal during the Cosmic Dawn has an amplitude on the order of 100 mK, the current system—at least from a thermal noise perspective—is sufficiently sensitive. For comparison, the HERA experiment used a similar integration time of 94 nights to place stringent upper limits on the 21 cm power spectrum during reionization, reaching sensitivities sufficient to rule out a broad class of cold IGM models (thou).
@@ 0.5 Future Work
We are developing novel, software optimized antenna designs (restrepo) along with radio sky model validation of antenna chromaticity and bayesian foreground models (Mora et al, in prep).
@@ 0.6 Conclusions
Compared to established radiometers such as EDGES, SARAS, PRIZM, and BIGHORNS, this CANTAR prototype demonstrates broadly consistent performance in key areas such as impedance match and system temperature, while introducing a novel digital backend architecture based on commercial software-defined radios (SDRs). For instance, our best-performing antenna configurations achieve simulated and measured $S_{11}$ values below –25 dB over bandwidths of 50–70 MHz, which is comparable to the –15 to –20 dB targets reported for EDGES high-band (bowman2009) and SARAS 3 (Singh_2018). Similarly, the prototype system temperature of 443–453 K using a dual-stage low-noise amplifier is within the 300–500 K range reported for BIGHORNS and PRIZM (Sokolowski_2015; prizm).